Q4: You want to buy a car, and a bank will lend you $30,000. The loan...

  • Q4: You want to buy a car, and a bank will lend you $30,000. The loan would be fully amortized over 3 years (36 months), and the nominal interest rate would be 5%, with interest paid monthly. What is the monthly loan payment? What is the loan's EFF%?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text General aCopy ,_a. ars-函Merge & Center, $, % , 弼,8 C Paste B l u. Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard Font Alignment Number Styles Cells FG97 FD FE FG FH FI FJ FK FL FM 83 84 85 86 87 LOAN RATE PERIODS IN A YEAR YEARS 30000 5% 12 89 90 91 92 93 94 95 96 97 98 ANS 1 PMT $ 899.13 EXCEL PMT(FH85/FH86,FH87 FH86,-FH84)] ANS 2 EFF 5.12% [FORMULA (1+0.05/12)A12-1] [FORMULA = (1+r/m)9m-1] m=12 100 101 CALCULATOR | LOAN OPTIONS fv. ann CAP STRU VALUE LOCK BOXDILUTION DOLLAR COST AVGMORTGAGE EXPO Sheeti Sheet2She 福 130% 07:07